# FiRST: Finetuning Router-Selective Transformers for Input-Adaptive Latency Reduction

### v2

@@ Abstract
Auto-regressive Large Language Models (LLMs) demonstrate remarkable performance across different domains such as vision and language processing. However, due to sequential processing through a stack of transformer layers, autoregressive decoding faces significant computation/latency challenges, particularly in resource-constrained environments like mobile and edge devices. Existing approaches in literature that aim to improve latency via skipping layers have two distinct flavors - 1) Early exit, and 2) Input-agnostic heuristics where tokens exit at pre-determined layers irrespective of input sequence. Both the above strategies have limitations - the former cannot be applied to handle KV Caching necessary for speed-ups in modern framework and the latter does not capture the variation in layer importance across tasks or more generally, across input sequences. To address both limitations, we propose FiRST, an algorithm that reduces inference latency by using layer-specific routers to select a subset of transformer layers adaptively for each input sequence - the prompt (during the prefill stage) decides which layers will be skipped during decoding. FiRST preserves compatibility with KV caching enabling faster inference while being quality-aware. FiRST is model-agnostic and can be easily enabled on any pre-trained LLM. Our approach reveals that input adaptivity is critical - indeed, different task-specific middle layers play a crucial role in evolving hidden representations depending on tasks. Extensive experiments show that FiRST significantly reduces latency while outperforming other layer selection strategies in quality metics. It retains competitive performance to base model (without layer skipping) and in some cases, even improves upon it. FiRST is thus a promising and efficient solution for LLM deployment in low-resource environments.
@@ 1 Introduction
Large Language Models (LLM’s) have revolutionized the fields of Natural Language Processing and Computer Vision achieving incredible performance on a diverse set of benchmark tasks. However, the scale of these LLM’s characterized by billions of parameters hinder their adoption in resource-constrained environments with memory, latency and compute being the main challenges. In this work, we focus on the latency aspect which becomes the most significant bottleneck for tasks such as machine translation, question answering, summarization particularly on devices, such as laptops and mobile phones. As noted by (Schuster et al., 2022), the auto-regressive nature of decoding in LLM’s further pronounces the latency bottleneck.
@@
Transformer based LLMs have several stacks of layers (including attention and FFN layers) leading to high latency and compute requirements, making inference very slow or even infeasible in resource constrained settings. This is because of the sequential processing of tokens through all the layers for every input sequence and task. However, it is important to note that in the real world, there is a lot of heterogeneity in input sequences and tasks. (Schuster et al., 2022; Sun et al., 2022) noted that the generations made by LLMs can have varying levels of difficulty and certain generations can be solved with reduced compute, by exiting the transformer stack early. At the same time, it has been noted in recent works (Wendler et al., 2024) that inference forward pass proceeds in phases through the layers of transformer based models, with different types of information being extracted or mapped at different phases (sequences of layers) for certain tasks such as translation. Motivated by these and other related works, we hypothesize that different sequential combinations of layers are important for different input sequences and tasks. Learning the right sequential combination of layers can help reduce inference latency and compute for on-device scenarios. However, there are several challenges. Any algorithm for determining the “right” combination of layers should minimize any quality loss, be compatible with other latency reduction strategies such as KV cache handling and be learnable with minimal compute/training overhead.
@@
In the last few years, several promising approaches have been proposed in literature that adaptively prune layers at each decoding step. Token-level early exit proposed in (Schuster et al., 2022; Sun et al., 2022) allow tokens to exit the transformer layer stack early based on different strategies to compute the confidence or saturation level. (Elhoushi et al., 2024; Elbayad et al., 2020; Zhang et al., 2019) extended this idea to incorporate layer skipping at a token level during training. While token level early exit is a useful idea in theory, it suffers from a major limitation of incompatible KV caching in practice (Del Corro et al., 2023). The incompatibility stems from having to recompute KV caches for preceding tokens if we have a delayed exit point for latter tokens, often resulting in loss of early exit advantages. This limits its practical adoption since KV cache is crucial in significantly speeding up auto-regressive decoding.
@@
Recently, (Liu et al., 2024; Del Corro et al., 2023; Song et al., 2024) have proposed input-agnostic layer skipping at token level, that handle KV cache appropriately as well as retain the advantage of adaptive partial computation. In these solutions, tokens exit at pre-determined layers irrespective of the input sequence, and for all sequences in a batch, tokens at the same position in a sequence exit at the same layer. Furthermore, tokens at latter parts of the sequence are constrained to exit earlier than the previous tokens to ensure that there is no redundant KV cache re-computation. These solutions are heuristic based and impose hard rules and constraints irrespective of input sequences, which can lead to drop in output quality. Others (Jaiswal et al., 2024; Chen et al., 2024) have proposed skipping layers by identifying redundant ones through computing cosine similarity of (input/output) representations of a layer. However, their strategy does not take into account that several middle layers are crucial (see (Liu et al., 2024)) and furthermore, final prediction capability of full model is not taken into account while deciding which layers to skip. Importantly, in none of the works described above, the strategy of selecting layers for skipping is sequence dependent. Furthermore, they do not consider finetuning the models in a way such that not only the performance improves but also the model learns to skip layers appropriately.
@@
Our goal is to design an input-adaptive learnable layer selection strategy with quality aware latency gains that is also able to handle the KV cache appropriately. Ideally, for every input sequence and task, we want to predict the optimal (sequential) combination of layers at inference time, such that quality loss is minimum and the latency gains are as high as possible. We want to do this with expending very little compute/additional training and appropriate handling of KV cache. We propose to do this via training routers. Based on the output of each layer for a sequence, a router will decide whether or not to skip the subsequent layer in the transformer architecture. Since the decision is at a sequence level, KV cache issues do not arise, as all tokens in a sequence would pass through the same set of layers. Finally, we fine-tune the model combined with trained routers using LoRA adapters to improve the quality significantly while retaining the latency gains. As an added bonus, LoRA finetuning smoothens the layer skipping and further highlights the varied importance of layers based on input sequence. We summarize our contributions:
@@
    We propose a training and inference algorithm FiRST that incorporates layer-specific routers for selecting layers in an input-adaptive manner. The layer selection is uniform for all tokens in a sequence, thus handling KV caches without introducing additional compute/latency. FiRST can be applied on top of any pre-trained model.
@@
    We propose LoRA based finetuning on top of router based layer selection to improve quality while retaining latency gains. This also smoothens layer selection.
@@
    Finally, we demonstrate an extensive set of experiments with FiRST on multiple datasets for 3 distinct tasks namely Machine Translation, Summarization and Question Answering, and 2 different open model architectures namely LLaMA-3-8B and LLaMA-3.2-3B. We show that for the same target speed-up, FiRST significantly improves performance across tasks as compared to baselines and prior works.
@@ 2 Problem Statement
Let us consider an an input sequence $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ with $n$ tokens. Let there be $m$ transformer layers in the model, where the $i^{th}$ transformer layer is represented as the function $\phi_{i}()$. As stated lucidly in (Wendler et al., 2024), $\mathcal{X}$ is first converted to an initial latent representation $\mathcal{H}_{0}=\{H^{1}_{0},H^{2}_{0},\ldots,H^{n}_{0}\}$, where $H^{0}_{j}\in\mathbb{R}^{D},\forall j\in[n]$ is a look-up from a learned embedding dictionary corresponding to the $j^{th}$ token. Thereafter, every transformer layer $\phi_{i}()$ operates on the latent vectors $\mathcal{H}_{i}$ to generate the embedding for the $i^{th}$ layer as follows. For the $j^{th}$ token,
@@ 4 FiRST Framework and Algorithm
In this section, we describe the training and inference frameworks for FiRST in details. We discuss how to train Routers to be adaptive to input sequences. Given an off-the-shelf pre-trained LLM, we propose two training phases. In the first phase, we train a router for each layer that decides whether the input sequence should skip the layer. In the second phase, to tackle the issue of unseen skipping during pre-training, we fine-tune the router-augmented LLM keeping router weights fixed to ensure the model improves performance on the target dataset without reducing the skipping level.
@@ 4.2 LoRA Compensation Module
Skip (%) Model Type English-to-German English-to-Chinese
BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 41.78 21.74 93 56.94 35.56 82.66
Base 37.17 18.57 87.13 38.02 22.46 68.95
15 Skip Decode Router + LoRA 23.04 10.52 55.62 28.73 15.84 55.98
Router 3.99 1.18 23.33 4.74 2.33 21.75
Random Skip Router + LoRA 30.43 10.98 66.25 47.88 25.75 67.32
Router 26.54 8.77 60.27 36.60 18.66 59.89
Unified Skip Router + LoRA 28.92 10.64 59.34 46.61 25.01 69.58
Router 23.23 7.85 59.26 27.28 13.35 54.57
FiRST (Ours) Router + LoRA 38.01 17.89 82.14 48.35 26.57 68.63
Router 28.83 11.8 67.74 17.55 8.68 42.76
25 Skip Decode Router + LoRA 13.67 6.00 31.47 20.03 10.85 33.85
Router 3.24 0.92 21.55 3.78 1.84 20.93
Random Skip Router + LoRA 6.01 0.91 29.71 11.69 4.66 27.73
Router 3.65 0.49 29.95 7.37 2.81 35.16
Unified Skip Router + LoRA 15.67 3.36 31.69 34.90 15.75 50.59
Router 12.58 2.65 32.15 17.74 7.35 38.74
FiRST (Ours) Router + LoRA 17.84 4.14 34.95 35.79 15.66 56.92
Router 9.67 1.37 26.01 11.01 3.23 25.45
Table 1: Machine Translation Results for LLaMA-3-8B: BLEU-1, BLEU-2, and COMET scores are reported for English-to-German and English-to-Chinese tasks across varying skip levels. FiRST consistently achieves the highest BLEU-1 and BLEU-2 at 15% skipping, for both translation directions. Similarly, high BLEU-1 and COMET scores are obtained for 25% skipping rate.
@@ 5 Experiments
Skip (%) Model Type English-to-German English-to-Chinese
BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 37.57 17.43 89.72 51.81 30.04 79.13
Base 31.19 13.67 81.66 32.10 17.92 61.84
15 Skip Decode Router + LoRA 23.24 11.07 44.58 38.14 21.68 46.70
Router 9.93 3.89 32.74 9.59 4.84 34.14
Random Skip Router + LoRA 18.99 4.78 47.26 38.75 17.39 57.79
Router 12.11 2.77 36.30 13.56 5.64 35.86
Unified Skip Router + LoRA 21.65 5.93 44.72 36.96 17.35 57.10
Router 16.41 3.99 39.81 22.43 9.37 45.16
FiRST (Ours) Router + LoRA 24.68 7.13 60.29 45.66 23.66 67.45
Router 16.47 4.11 43.04 22.69 10.92 54.55
25 Skip Decode Router + LoRA 16.85 7.62 32.33 27.81 15.74 42.01
Router 9.14 3.51 27.64 7.01 3.41 29.51
Random Skip Router + LoRA 8.95 1.07 30.97 25.12 9.85 45.53
Router 5.32 0.69 27.22 10.07 3.69 31.50
Unified Skip Router + LoRA 15.14 2.65 39.81 30.56 12.27 42.30
Router 10.21 1.62 30.86 15.21 5.51 31.63
FiRST (Ours) Router + LoRA 18.89 4.06 45.38 32.92 13.74 41.66
Router 9.51 1.44 29.08 10.46 3.55 27.83
Table 2: Machine Translation Results for LLaMA-3.2-3B: BLEU-1, BLEU-2, and COMET scores are shown for English-to-German and English-to-Chinese tasks at different skip levels. FiRST consistently achieves the highest BLEU-1 and COMET scores for English-to-German at both skipping percentages and highest BLEU-1 for English-to-Chinese.
Skip (%) Model Type EM F1
0 Original Model Base + LoRA 73.93 85.99
Base 19.46 36.73
10 Skip Decode R + LoRA 60.14 65.33
Router 16.38 31.48
Random Skip R + LoRA 65.73 80.08
Router 18.25 33.75
Unified Skip R + LoRA 55.54 74.58
Router 17.39 32.91
FiRST (Ours) R + LoRA 70.85 83.61
Router 14.58 31.52
20 Skip Decode R + LoRA 45.00 55.10
Router 10.68 26.69
Random Skip R + LoRA 47.79 66.37
Router 6.71 22.46
Unified Skip R + LoRA 52.87 69.28
Router 18.18 32.51
FiRST (Ours) R + LoRA 60.60 75.49
Router 13.21 27.48
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.07 84.17
Base 18.92 37.74
10 Skip Decode R + LoRA 60.79 75.00
Router 20.00 31.55
Random Skip R + LoRA 64.78 77.27
Router 13.76 28.59
Unified Skip R + LoRA 65.03 77.53
Router 13.16 32.31
FiRST R + LoRA 69.44 81.35
Router 12.79 28.37
20 Skip Decode R + LoRA 40.12 40.00
Router 20.45 37.62
Random Skip R + LoRA 11.32 38.34
Router 6.75 15.51
Unified Skip R + LoRA 37.39 52.49
Router 7.81 18.20
FiRST R + LoRA 39.70 54.59
Router 5.52 15.33
Table 3: Quality Analysis on Question Answering (SQuAD v1.1): LLaMA-3-8B (left) and LLaMA-3.2-3B (right). Exact Match (EM) and F1 scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
@@
0 Original Model wLoRA 84.87 28.46 16.99
@@
27 Skip Decode R + LoRA 79.92 10.67 10.32
Router 77.27 9.59 7.00
Random Skip R + LoRA 76.40 11.45 7.89
Router 77.45 12.56 9.08
Unified Skip R + LoRA 80.28 15.94 9.89
Router 77.43 10.97 7.68
FiRST (Ours) R + LoRA 77.5 14.65 10.45
Router 75.6 9.39 6.92
@@
0 Original Model wLoRA 84.89 28.37 17.02
@@
Table 4: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
We conduct experiments on three benchmark tasks namely Machine Translation, Text Summarization and Question Answering demonstrating both robustness and scalability of FiRST.
Datasets: For machine translation, we use WMT development sets (2017–2020) for English-to-Chinese and English-to-German tasks, evaluating performance on the WMT 2022 test set, which covers diverse domains such as news, social media, e-commerce, and conversational contexts. For summarization, we use the CNN/DailyMail dataset, with 4000 randomly selected training samples and evaluation on the standard test set of 11,490 samples. In Question Answering, we utilize SQuAD v1.0, training on 4000 randomly selected samples from the 100k question-answer pairs and evaluating on the validation set, as test set labels are unavailable. Appendix A.2 contains detailed descriptions.
Evaluation Metrics: We use standard metrics to evaluate the quality of generated output in each of the three tasks. For Machine Translation, we benchmark using BLEU scores and COMET where the latter provides a more nuanced assessment beyond $n$-gram used in BLEU. For Summarization, we use ROUGE scores and BERT Score - as before, the latter captures meaning-based similarity. For Question Answering, we use Exact Match and F1 Score as the metrics to benchmark the output quality. Finally, for benchmarking latency, we look at the TPOT (Time Per Output Token). TPOT evaluates the average time taken to produce each output token and is calculated for GPU to gauge overall decoding performance. Appendix A.3 contains a detailed description of evaluation metrics. Additionally, hyper-parameters used during training and inference can be found in Appendix A.4.

### v3

@@ Abstract
Auto-regressive Large Language Models (LLMs) demonstrate remarkable performance across different domains such as vision and language tasks. However, due to sequential processing through multiple transformer layers, autoregressive decoding faces significant computational challenges, particularly in resource-constrained environments like mobile and edge devices. Existing approaches in literature that aim to improve latency via skipping layers have two distinct flavors: (1) early exit, and (2) input-agnostic heuristics where tokens exit at pre-determined layers irrespective of input sequence. Both the above strategies have limitations—the former cannot be applied in the presence of KV caching, which is essential for speed-ups in modern inference frameworks, and the latter fails to capture variation in layer importance across tasks or, more generally, across input sequences. To address these limitations, we propose FiRST, a model-agnostic framework that reduces inference latency by using layer-specific routers to adaptively skip transformer layers during decoding, based on routing decisions made from the input prompt in the prefill stage. FiRST remains fully compatible with KV caching, enabling faster decoding while maintaining quality. Our method reveals that input adaptivity is essential—different tasks rely on different subsets of layers to evolve meaningful representations. Extensive experiments show that FiRST significantly reduces latency while outperforming existing layer selection strategies in quality. It retains performance comparable to the base model without skipping. FiRST is thus a promising and efficient solution for LLM deployment in low-resource environments.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized the fields of Natural Language Processing and Computer Vision achieving incredible performance on a diverse set of benchmark tasks. However, the massive scale of LLMs—often involving billions of parameters—poses significant challenges for deployment in resource-constrained environments, where memory, compute, and especially latency become critical bottlenecks. In this work, we focus on addressing the latency issue, which is particularly pronounced in edge settings such as laptops and mobile devices. As noted by Schuster et al. (2022a), the auto-regressive nature of decoding in LLMs further amplifies this bottleneck.
@@
Transformer-based LLMs consist of multiple stacked layers—including attention and feed-forward networks—which result in high latency and computational cost. This makes inference slow or even impractical in resource-constrained environments. The inefficiency stems from the need to process each input token sequentially through all layers, irrespective of the input sequence or task. However, it is important to note that in the real world, there is a lot of heterogeneity in input sequences and tasks. (Schuster et al., 2022a; Sun et al., 2022) observed that LLM generations can have varying levels of difficulty and certain generations can be solved with reduced computation, by exiting the transformer stack early. At the same time, it has been noted in recent works (Wendler et al., 2024) that inference forward pass proceeds in phases through the layers of transformer-based models, with different types of information being extracted or mapped at different phases (sequences of layers) for certain tasks such as translation. Motivated by these and other related works, we hypothesize that different sequential combinations of layers are important for different input sequences and tasks. Learning the right sequential combination of layers can help reduce inference latency and compute for on-device scenarios. However, there are several challenges. Any algorithm for determining the “right” combination of layers should minimize any quality loss, be compatible with other latency reduction strategies such as KV cache handling, and be learnable with minimal compute/training overhead.
@@
In the last few years, several promising approaches have been proposed in literature that adaptively prune layers at each decoding step. Token-level early exit proposed in (Schuster et al., 2022a; Sun et al., 2022) allow tokens to exit the transformer layer stack early based on different strategies to compute the confidence or saturation level. (Elhoushi et al., 2024; Elbayad et al., 2020; Zhang et al., 2019) extended this idea to incorporate layer skipping at a token level during training. While token level early exit is a useful idea in theory, it suffers from a major limitation of incompatible KV caching in practice (Del Corro et al., 2023). The incompatibility stems from having to recompute KV caches for preceding tokens if we have a delayed exit point for latter tokens, often resulting in loss of early exit advantages. This limits its practical adoption since KV cache is crucial in significantly speeding up auto-regressive decoding.
@@
Recently, (Liu et al., 2024; Del Corro et al., 2023; Song et al., 2024) have proposed input-agnostic layer skipping at token level, that handle KV cache appropriately as well as retain the advantage of adaptive partial computation. In these solutions, tokens exit at pre-determined layers irrespective of the input sequence, and for all sequences in a batch, tokens at the same position in a sequence exit at the same layer. Furthermore, tokens at latter parts of the sequence are constrained to exit earlier than the previous tokens to ensure that there is no redundant KV cache re-computation. These solutions are heuristic based and impose hard rules and constraints irrespective of input sequences, which can lead to drop in output quality. Others (Jaiswal et al., 2024; Chen et al., 2024) have proposed skipping layers by identifying redundant ones through computing cosine similarity of (input/output) representations of a layer. However, their strategy does not take into account that several middle layers are crucial (see (Liu et al., 2024)) and furthermore, final prediction capability of full model is not taken into account while deciding which layers to skip. Importantly, in none of the works described above, the strategy of selecting layers for skipping is sequence dependent. Furthermore, they do not consider fine-tuning the models in a way such that not only the performance improves but the model also learns to skip layers appropriately.
@@
Our goal is to design an input-adaptive, learnable layer selection strategy that provides quality-aware latency improvements while properly handling KV caching. Ideally, for each input sequence and task, we want to predict a sequential combination of layers to run during inference—minimizing quality loss while achieving as much latency gain as possible. We also want to do this with minimal computational overhead or extra training. To achieve this, we propose training routers. At each layer, the router looks at the current sequence representation and decides whether to skip the next layer. Since the decision is made at the sequence level, all tokens follow the same path through the model, avoiding KV cache inconsistencies during decoding.Finally, we fine-tune the model with trained routers using LoRA adapters to recover any quality drop introduced by layer skipping, while preserving latency gains. LoRA fine-tuning also smoothens layer skipping and further highlights the varied importance of layers based on input sequence. Our key contributions include:
@@
    We propose FiRST, a training and inference algorithm that uses layer-specific routers to perform input-adaptive layer selection. All tokens in a sequence follow the same selected layers, ensuring compatibility with KV caching and avoiding additional compute or latency overhead. FiRST is model-agnostic and can be applied on top of any pre-trained LLM.
@@
    We introduce a LoRA-based fine-tuning approach on top of router-based layer selection to recover quality while maintaining latency gains. This also encourages smoother and a more stable layer selection.
@@
    Finally, we conduct extensive experiments with FiRST across multiple datasets spanning three distinct tasks: Machine Translation, Summarization, and Question Answering. We evaluate on two open-source model architectures—LLaMA-3-8B and LLaMA-3.2-3B—and show that, for the same target speed-up, FiRST significantly improves performance across tasks as compared to baselines.
@@ 2 Problem Statement
Let us consider an an input sequence $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ with $n$ tokens. Let there be $m$ transformer layers in the model, where the $i^{th}$ transformer layer is represented as the function $\phi_{i}()$. As stated lucidly in (Wendler et al., 2024), $\mathcal{X}$ is first converted to an initial latent representation $\mathcal{H}_{0}=\{H^{1}_{0},H^{2}_{0},\ldots,H^{n}_{0}\}$, where $H^{0}_{j}\in\mathbb{R}^{D},\forall j\in[n]$ is a look-up from a learned embedding dictionary corresponding to the $j^{th}$ token. Thereafter, every transformer layer $\phi_{i}()$ operates on the latent vectors $\mathcal{H}_{i}$ to generate the embedding for the $i^{th}$ layer as follows. For the $j^{th}$ token, the embedding at layer $i$ is computed as follows:
@@ 4 FiRST Framework and Algorithm
In this section, we describe the training and inference frameworks for FiRST in details. We discuss how to train routers to be adaptive to input sequences. Given an off-the-shelf pre-trained LLM, we propose two training phases. In the first phase, we train a router for each layer that decides whether the input sequence should skip the layer. In the second phase, to tackle the issue of unseen skipping during pre-training, we fine-tune the router-augmented LLM keeping router weights fixed to ensure the model improves performance on the target dataset without reducing the skipping level. While joint training of the router and LoRA modules is theoretically possible, we find it introduces optimization instability (see Appendix A.4 for further discussion on this design choice and its impact).
@@ 4.2 LoRA Compensation Module
En-to-De En-to-Zh
Skip (%) Model Type LLaMA-3-8B LLaMA-3.2-3B LLaMA-3-8B LLaMA-3.2-3B
BLEU COMET BLEU COMET BLEU COMET BLEU COMET
0 Original Model Base + LoRA 0.199 93.00 0.160 89.72 0.333 82.66 0.278 79.13
Base 0.169 87.13 0.125 81.66 0.208 68.95 0.166 61.84
15 Skip Decode Router + LoRA 0.094 55.62 0.105 44.58 0.149 55.98 0.2 46.7
Router 0.019 23.33 0.055 32.74 0.03 21.75 0.062 34.14
Random Skip Router + LoRA 0.097 66.25 0.079 47.26 0.237 67.32 0.154 57.79
Router 0.132 60.27 0.048 36.3 0.168 59.89 0.077 35.86
Unified Skip Router + LoRA 0.153 59.34 0.095 44.72 0.23 69.58 0.157 57.1
Router 0.117 59.26 0.067 39.81 0.122 54.57 0.087 45.16
FiRST (Ours) Router + LoRA 0.161 82.14 0.113 60.29 0.247 68.63 0.218 67.45
Router 0.108 67.74 0.069 43.04 0.08 42.76 0.1 54.55
25 Skip Decode Router + LoRA 0.07 31.47 0.077 32.33 0.088 33.85 0.147 42.01
Router 0.015 21.55 0.05 27.64 0.024 20.93 0.045 29.51
Random Skip Router + LoRA 0.018 29.71 0.023 30.97 0.065 27.73 0.137 45.53
Router 0.011 29.95 0.015 27.22 0.04 35.16 0.053 31.5
Unified Skip Router + LoRA 0.06 31.69 0.05 39.81 0.142 50.59 0.108 42.3
Router 0.048 32.15 0.032 30.86 0.068 38.74 0.079 31.63
FiRST (Ours) Router + LoRA 0.071 34.95 0.072 45.38 0.119 56.92 0.126 41.66
Router 0.029 26.01 0.029 29.08 0.051 25.45 0.053 27.83
Table 1: Machine Translation results for the English-to-German and English-to-Chinese task for both models: BLEU and COMET scores are reported across varying skip levels. FiRST (Ours) performs consistently well across both translation directions.
@@ 5 Experiments
Skip (%) Model Type SQuAD NQ
EM F1 EM
LLaMA-3-8B
0 Original Model Base + LoRA 73.93 85.99 51.88
Base 19.46 36.73 37.40
10 Skip Decode R + LoRA 60.14 65.33 41.98
Router 16.38 31.48 33.22
Random Skip R + LoRA 65.73 80.08 44.95
Router 18.25 33.75 34.09
Unified Skip R + LoRA 55.54 74.58 45.91
Router 17.39 32.91 33.64
FiRST (Ours) R + LoRA 70.85 83.61 47.85
Router 14.58 31.52 33.53
20 Skip Decode R + LoRA 45.00 55.10 26.62
Router 10.68 26.69 14.39
Random Skip R + LoRA 47.79 66.37 28.96
Router 6.71 22.46 27.40
Unified Skip R + LoRA 52.87 69.28 25.30
Router 18.18 32.51 23.57
FiRST (Ours) R + LoRA 60.60 75.49 32.22
Router 13.21 27.48 18.02
LLaMA-3.2-3B
0 Original Model Base + LoRA 73.07 84.17 40.50
Base 18.92 37.74 30.10
10 Skip Decode R + LoRA 60.79 75.00 31.74
Router 20.00 31.55 21.51
Random Skip R + LoRA 64.78 77.27 36.02
Router 13.76 28.59 22.76
Unified Skip R + LoRA 65.03 77.53 32.90
Router 13.16 32.31 21.03
FiRST (Ours) R + LoRA 69.44 81.35 37.82
Router 12.79 28.37 22.55
20 Skip Decode R + LoRA 40.12 40.00 27.28
Router 20.45 37.62 14.73
Random Skip R + LoRA 11.32 38.34 26.60
Router 6.75 15.51 13.24
Unified Skip R + LoRA 37.39 52.49 26.84
Router 7.81 18.20 16.09
FiRST (Ours) R + LoRA 39.70 54.59 29.87
Router 5.52 15.33 17.51
Table 2: Quality Analysis on Question Answering tasks using LLaMA-3-8B and LLaMA-3.2-3B across SQuAD (Exact Match and F1) and Natural Questions (Exact Match). Results are shown for multiple skip strategies and levels. FiRST (Ours) consistently performs best under skipping, demonstrating strong robustness. R + LoRA indicates Router Augmentation followed by LoRA fine-tuning.
@@
LLaMA-3-8B
0 Original Model Base + LoRA 84.87 28.46 16.99
@@
20 Skip Decode R + LoRA 82.57 20.41 14.87
Router 81.62 13.48 9.19
Random Skip R + LoRA 81.39 21.57 13.83
Router 79.23 15.51 10.93
Unified Skip R + LoRA 82.93 22.3 13.37
Router 80.32 16.51 11.15
FiRST (Ours) R + LoRA 82.8 27.65 17.84
Router 79.32 16.28 10.85
@@
LLaMA-3.2-3B
0 Original Model Base + LoRA 84.89 28.37 17.02
@@
Table 3: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework).
We conduct experiments on three benchmark tasks—Machine Translation, Text Summarization, and Question Answering—demonstrating the robustness and scalability of FiRST across diverse settings. We base our task selection on prior work in the field to ensure a fair and meaningful comparison Liu et al. (2024); Del Corro et al. (2023); Schuster et al. (2022b).
Datasets: For machine translation, we use WMT development sets (2017–2020) for English-to-Chinese and English-to-German tasks, evaluating performance on the WMT 2022 test set, which covers diverse domains such as news, social media, e-commerce, and conversational contexts. For summarization, we use the CNN/DailyMail dataset, with 4,000 randomly selected training samples and evaluation on the standard test set of 11,490 samples. For question answering (QA), we utilize SQuAD v1.1 and Natural Questions (NQ). We train on 4,000 randomly selected samples from each dataset and evaluate on their respective validation sets as test set labels are unavailable. For NQ, we incorporate a retrieval step before answer generation, retrieving relevant passages as context. Appendix A.2 contains detailed descriptions.
Evaluation Metrics: We employ standard metrics to assess output quality across three distinct tasks. For Machine Translation, we benchmark performance using BLEU and COMET; COMET is included for its more nuanced assessment capabilities beyond the n-gram overlap measured by BLEU. Summarization quality is evaluated with ROUGE scores and BERT Score, the latter capturing meaning-based similarity. For Question Answering, we use Exact Match (EM) and F1 Score on the SQuAD dataset, and report Exact Match (EM) for Natural Questions (NQ). Finally, to benchmark latency, we measure Time Per Output Token (TPOT) on GPU, which gauges overall decoding performance. Detailed descriptions of all evaluation metrics are available in Appendix A.3. Hyperparameters used during training and inference are provided in Appendix A.4.